# VideoPainter: Any-length Video Inpainting and Editing with Plug-and-Play Context Control

### v2

@@ Abstract
Video inpainting, crucial for the media industry, aims to restore corrupted content. However, current methods relying on limited pixel propagation or single-branch image inpainting architectures face challenges with generating fully masked objects, balancing background preservation with foreground generation, and maintaining ID consistency over long video. To address these issues, we propose VideoPainter, an efficient dual-branch framework featuring a lightweight context encoder. This plug-and-play encoder processes masked videos and injects background guidance into any pre-trained video diffusion transformer, generalizing across arbitrary mask types, enhancing background integration and foreground generation, and enabling user-customized control. We further introduce a strategy to resample inpainting regions for maintaining ID consistency in any-length video inpainting. Additionally, we develop a scalable dataset pipeline using advanced vision models and construct VPData and VPBench—the largest video inpainting dataset with segmentation masks and dense caption ($>390K$ clips) —to support large-scale training and evaluation. We also show VideoPainter’s promising potential in downstream applications such as video editing. Extensive experiments demonstrate VideoPainter’s state-of-the-art performance in any-length video inpainting and editing across $8$ key metrics, including video quality, mask region preservation, and textual coherence.
Machine Learning, ICML
Figure 1: VideoPainter enables plug-and-play text-guided video inpainting and editing for any video length and pre-trained Diffusion Transformer with masked video and video caption (user editing instruction). The upper part demonstrates the effectiveness of VideoPainter in various video inpainting scenarios, including object, landscape, human, animal, multi-region (Multi), and random masks. The lower section demonstrates the performance of VideoPainter in video editing, including adding, removing, changing attributes, and swapping objects. In both video inpainting and editing, we demonstrate strong ID consistency in generating long videos (Any Len.).
1 Introduction
@@
Video inpainting (Quan et al., 2024), which aims to restore the corrupted video while maintaining coherence, facilitates numerous applications, including try-on (Fang et al., 2024), film production (Polyak et al., 2024), and video editing (Sun et al., 2024). Recently, Diffusion Transformers (DiT) (Peebles & Xie, 2023; OpenAI, 2024) have shown promise in video generation, leading to the exploration of generative video inpainting (Zhang et al., 2024b; Zi et al., 2024).
@@
Figure 2: Framework Comparison. Non-generative approaches, limited to pixel or pixel feature propagation from backgrounds, fail to inpaint fully segmentation-masked objects due to the lack of enough background context infomation. Generative methods adapt single-branch image inpainting models to video by adding temporal attention, struggling to maintain background fidelity and generate foreground contents in one model. In contrast, VideoPainter implements a dual-branch architecture that leverages an efficient context encoder with any pre-trained DiT, decoupling video inpainting to background preservation and foreground generation, and enabling plug-and-play video inpainting control.
@@
Existing approaches, as illustrated in Fig. 2, can be broadly categorized into two types: (1) Non-Generative methods (Zhou et al., 2023; Li et al., 2022; Lee et al., 2019) depend on limited pixel feature propagation (physical constraints or model architectural priors), which only take masked videos as inputs and cannot generate fully segmentation-masked objects. (2) Generative methods (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024) extend single-branch image inpainting architectures (Rombach & Esser, 2022) to video by incorporating temporal attention, which struggles to balance background preservation and foreground generation in one model and obtain inferior temporal coherence compared to native video DiTs. Moreover, both paradigms neglect long video inpainting and struggle to maintain consistent object ID with long videos.
@@
This motivates us to decompose video inpainting into background preservation and foreground generation and adopt a dual-branch architecture in DiTs, where we can incorporate a dedicated context encoder for masked video feature extraction while utilizing the pre-trained DiT’s capabilities to generate semantic coherent video content conditioned on both the preserved background and text prompts. Similar observations have been made in image inpainting research, notably in BrushNet (Ju et al., 2024) and ControlNet (Zhang et al., 2023). However, directly applying their architecture to video DiTs presents several challenges: (1) Given Video DiT’s robust generative foundation and heavy model size, replicating the full/half-giant Video DiT backbone as the context encoder would be unnecessary and computationally prohibitive. (2) Unlike BrushNet’s pure convolutional control branch, DiT’s tokens in masked regions inherently contain background information due to global attention, complicating the distinction between masked and unmasked regions in DiT backbones. (3) ControlNet lacks dense feature injection across all backbone layers, hindering dense background control for inpainting tasks.
@@
To address these challenges, we introduce VideoPainter, which enhances pre-trained DiT with a lightweight context encoder comprising only 6% of the backbone parameters, to form the first efficient dual-branch video inpainting architecture. VideoPainter features three main components: (1) A streamlined context encoder with just two layers, which integrates context features into the pre-trained DiT in a group-wise manner, ensuring efficient and dense background guidance. (2) Mask-selective feature integration to clearly distinguish the tokens of the masked and unmasked region. (3) A novel inpainting region ID resampling technique to efficiently process videos of any length while maintaining ID coherence. By freezing the pre-trained context encoder and DiT backbone and adding an ID-Adapter, we enhance the backbone’s attention sampling by concatenating the original key-value vectors with the inpainting region tokens. During inference, inpainting region tokens from previous clips are appended to the current KV vectors, ensuring the long-term preservation of target IDs. Notably, our VideoPainter supports plug-and-play and user-customized control.
@@
  • We propose VideoPainter, the first dual-branch video inpainting diffusion transformer framework that supports plug-and-play and user-customized control.
  • We design a lightweight context encoder for efficient and dense background control, and an inpainting region ID resampling for boosting the target object ID consistency in any-length video inpainting and editing.
@@
  • Experiments show VideoPainter achieves state-of-the-art performance across $8$ key metrics, including video quality, masked region preservation, and text alignment in video inpainting and video editing.
2 Related Work
2.1 Video Inpainting
Video inpainting approaches can be broadly classified into two categories based on whether they are generative:
@@ 2.1 Video Inpainting
These methods (Zhou et al., 2023; Li et al., 2022; Hu et al., 2020; Zhang et al., 2022c, b) leverage architecture priors to facilitate pixel propagation. This includes utilizing local perception of 3D CNNs (Chang et al., 2019a; Wang et al., 2019; Hu et al., 2020; Chang et al., 2019b), and exploiting the global perception of attention to retrieve and aggregate tokens with similar texture for filling masked video (Lee et al., 2019; Zeng et al., 2020; Liu et al., 2021; Zhang et al., 2022c). They also introduce various physical quantities, especially optical flow, as auxiliary conditions as it simplifies RGB pixel inpainting by completing less complex flow fields (Kim et al., 2019; Li et al., 2020; Zou et al., 2021; Xu et al., 2019; Gao et al., 2020; Zhang et al., 2022a, b). However, they are only effective for partial object occlusions with random masks but face significant limitations when inpaint fully masked regions due to insufficient contexts.
@@
Recent advances in generative foundation models (Rombach et al., 2022; Guo et al., 2023) have sparked numerous approaches that leverage additional modules or training strategies to extend backbones’ capabilities for video inpainting (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024). AVID (Zhang et al., 2024b) and COCOCO (Zi et al., 2024) represent the most related recent works. Both adopt a similar implementation by augmenting Stable Diffusion Inpainting (Rombach & Esser, 2022) with trainable temporal attention layers. This architecture includes per-frame region filling based on the image inpainting backbone and temporal smoothing with temporal attention. Despite showing promising results for both random and segmentation masks due to their generative abilities, they struggle to balance background preservation and foreground generation with text caption (Ju et al., 2024; Li et al., 2024) within the single backbone. AVID also explores any-length video inpainting by smoothing latent at segment boundaries and using the middle frame as the ID reference. In contrast, VideoPainter is a dual-branch framework by decoupling video inpainting into foreground generation and background-guided preservation. It employs an efficient context encoder to guide any pre-trained DiT, facilitating plug-and-play control. Furthermore, VideoPainter also introduces a novel inpainting region ID resampling technique that enables ID consistency in any-length video inpainting.
Table 1: Comparison of video inpainting datasets. Our VPData is the largest video inpainting dataset to date, comprising over 390K high-quality clips with segmentation masks, video captions, and masked region descriptions.
@@
2.2 Video Inpainting Datasets
Recent advances in segmentation (Ravi et al., 2024) have created many video segmentation datasets (Perazzi et al., 2016; Xu et al., 2018; Hong et al., 2023; Ding et al., 2023; Tokmakov et al., 2023; Darkhalil et al., 2022). Among these, DAVIS (Perazzi et al., 2016) and YouTube-VOS (Xu et al., 2018) have become prominent benchmarks for video inpainting due to their high-quality masks and diverse object categories. However, the existing datasets face two primary limitations: (1) insufficient scale to meet the data requirements of generative models, and (2) the absence of crucial control conditions necessary for generating masked objects such as video captions. In contrast, as shown in Tab. 1, we developed a scalable dataset pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), and constructed the largest video inpainting dataset to date with over 390K clips, each annotated with precise segmentation mask sequences and high-quality dense video captions.
Figure 3: Dataset Construction Pipeline. It consists of five dataset pre-processing steps: collection, annotation, splitting, selection, and captioning.
3 Method
@@
Figure 4: Model overview. The upper figure shows the architecture of VideoPainter. The context encoder performs video inpainting based on concatenation of the noisy latent, downsampled masks, and masked video latent via VAE. Features extracted by the context encoder are integrated into the pre-trained DiT in a group-wise and token-selective manner, where two encoder layers modulate the first and second halves of the DiT, respectively, and only the background tokens will be integrated into the backbone to prevent information ambiguity. The lower figure illustrates the inpainting ID region resampling with the ID Resample Adapter. During training, tokens of the current masked region are concatenated to the KV vectors, enhancing ID preservation of the inpainting region. During inference, the ID tokens of the last clip are concatenated to the current KV vectors, maintaining ID consistency with the last clip by resampling.
3.1 VPData and VPBench Construction Pipeline
@@
Collection. We chose Videvo and Pexels  [^1] Videvo: https://www.videvo.net/, Pexels: https://www.pexels.com/ as our data sources. We finally obtained around $450K$ publicly available internet videos from these sources.
@@
  • We employ the Recognize Anything Model (Zhang et al., 2024a) for open-set video tagging to identify the primary objects in the given videos.
  • Based on the detected primary object tags, we utilize Grounding DINO (Liu et al., 2023) to detect the corresponding bounding boxes for objects at fixed intervals.
  • These bounding boxes serve as prompts for SAM2 (Ravi et al., 2024), which generates precise mask segmentations.
Splitting. Scene transitions may occur while tracking the same object from different angles, causing disruptive view changes. We utilize PySceneDetect (Castellano, 2024) to identify scene transitions and subsequently partition the masks. Then we segmented the sequences into 10-second intervals and discarded short clips ($<6s$).
Selection. We employ $3$ key criteria: (1) Aesthetic Quality, evaluated using the Laion-Aesthetic Score Predictor (Schuhmann et al., 2022); (2) Motion Strength, predicted by optical flow measurements using the RAFT(Teed & Deng, 2020); and (3) Content Safety, assessed via the Stable Diffusion Safety Checker (Rombach et al., 2022).
@@
3.2 Dual-branch Inpainting Control
We incorporate masked video features into the pre-trained diffusion transformer (DiT) via an efficient context encoder, to decouple the background context extraction and foreground generation. This encoder processes a concatenated input of noisy latent, masked video latent, and downsampled masks. Specifically, the noisy latent provides information about the current generation. The masked video latent, extracted via VAE, aligns with the pre-trained DiT’s latent distribution. We apply cubic interpolation to down-sample masks, ensuring dimensional compatibility.
@@ 3.2 Dual-branch Inpainting Control
3.3 Target Region ID Resampling
While current video DiTs have shown great promise in handling temporal dynamics (Kuaishou, 2024; Bian et al., 2024), they often struggle to maintain smooth transitions and long-term identity consistency.
Smooth Transition. Following AVID (Zhang et al., 2024b), we employ overlapping generation and weighted average to maintain consistent transitions. Additionally, we utilize the last frame of the last generated video clip (before overlap) as the first frame of the current video clip’s overlapping region to ensure visual appearance continuity.
Identity Consistency. To maintain identity consistency in the long video, we introduce an inpainting region ID resampling method, as shown in lower Fig. 4. During training, we freeze both the DiT and the context encoder. Then we add trainable ID-Resample Adapters into the frozen DiT (LoRA), enabling ID resampling functionality. Specifically, tokens from the current masked region, which contain the desired ID, are concatenated with the KV vectors, thereby enhancing ID preservation in the inpainting region through additional KV resampling. Specifically, given current $\mathbf{Q}^{v}_{i}$, $\mathbf{K}^{v}_{i}$, and $\mathbf{V}^{v}_{i}$, we filter the mask region tokens in current $\mathbf{K}^{v}_{i}$ and $\mathbf{V}^{v}_{i}$, and concatenate them to $\mathbf{K}^{v}_{i}$ and $\mathbf{V}^{v}_{i}$, forcing the model to resample these tokens that have the needed ID. During inference, we prioritize maintaining ID consistency with the inpainting region tokens from the previous clip, as it represents the most temporally proximate generated result. Therefore, we concatenate the masked region tokens from the previous clip with the current key-value vectors, effectively resampling and maintaining the identity information in long video processing.
3.4 Plug-and-Play Control

### v3

@@ Abstract.
Video inpainting, crucial for the media industry, aims to restore corrupted content. However, current methods relying on limited pixel propagation or single-branch image inpainting architectures face challenges with generating fully masked objects, balancing background preservation with foreground generation, and maintaining ID consistency over long video. To address these issues, we propose VideoPainter, an efficient dual-branch framework featuring a lightweight context encoder. This plug-and-play encoder processes masked videos and injects background guidance into any pre-trained video diffusion transformer, generalizing across arbitrary mask types, enhancing background integration and foreground generation, and enabling user-customized control. We further introduce a strategy to resample inpainting regions for maintaining ID consistency in any-length video inpainting. Additionally, we develop a scalable dataset pipeline using advanced vision models and construct VPData and VPBench—the largest video inpainting dataset with segmentation masks and dense caption (¿390K clips) —to support large-scale training and evaluation. We also show VideoPainter’s promising potential in downstream applications such as video editing. Extensive experiments demonstrate VideoPainter’s state-of-the-art performance in any-length video inpainting and editing across $8$ key metrics, including video quality, mask region preservation, and textual coherence.
Artificial Intelligence Generative Content, Computer Vision, Video Inpainting, Video Editing
submissionid: 626journal: TOGccs: Computing methodologies Computer vision
1. Introduction
@@
Video inpainting (Quan et al., 2024), which aims to restore the corrupted video while maintaining coherence, facilitates numerous applications, including try-on (Fang et al., 2024), film production (Polyak et al., 2024), and video editing (Sun et al., 2024). Recently, Diffusion Transformers (DiT) (Peebles and Xie, 2023; OpenAI, 2024) have shown promise in video generation, leading to the exploration of generative video inpainting (Zhang et al., 2024b; Zi et al., 2024).
@@
Figure 2. Framework Comparison. Non-generative approaches, limited to pixel propagation from backgrounds, fail to inpaint fully segmentation-masked objects. Generative methods adapt single-branch image inpainting models to video by adding temporal attention, struggling to maintain background fidelity and generate foreground contents in one model. In contrast, VideoPainter implements a dual-branch architecture that leverages an efficient context encoder with any pre-trained DiT, decoupling video inpainting to background preservation and foreground generation, and enabling plug-and-play video inpainting control.
@@
Existing approaches, as illustrated in Fig. 2, can be broadly categorized into two types: (1) Non-Generative methods (Zhou et al., 2023; Li et al., 2022; Lee et al., 2019) depend on limited pixel feature propagation (physical constraints or model architectural priors), which only take masked videos as inputs and cannot generate fully segmentation-masked objects. (2) Generative methods (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024) extend single-branch image inpainting architectures (Rombach and Esser, 2022) to video by incorporating temporal attention, which struggles to balance background preservation and foreground generation in one model and obtain inferior temporal coherence compared to native video DiTs. Moreover, both paradigms neglect long video inpainting and struggle to maintain consistent object identity with long videos.
@@
This motivates us to decompose video inpainting into background preservation and foreground generation and adopt a dual-branch architecture in DiTs, where we can incorporate a dedicated context encoder for masked video feature extraction while utilizing the pre-trained DiT’s capabilities to generate semantic coherent video content conditioned on both the preserved background and text prompts. Similar observations have been made in image inpainting research, notably in BrushNet (Ju et al., 2024) and ControlNet (Zhang et al., 2023). However, directly applying their architecture to video DiTs presents several challenges: (1) Given Video DiT’s robust generative foundation and heavy model size, replicating the full/half-giant Video DiT backbone as the context encoder would be unnecessary and computationally prohibitive. (2) Unlike BrushNet’s pure convolutional control branch, DiT’s tokens in masked regions inherently contain background information due to global attention, complicating the distinction between masked and unmasked regions in DiT backbones. (3) ControlNet lacks feature injection across all layers, hindering dense background control for inpainting tasks.
@@
To address these challenges, we introduce VideoPainter, which enhances pre-trained DiT with a lightweight context encoder comprising only 6% of the backbone parameters, to form the first efficient dual-branch video inpainting architecture. VideoPainter features three main components: (1) A streamlined context encoder with just two layers, which integrates context features into the pre-trained DiT in a group-wise manner, ensuring efficient and dense background guidance. (2) Mask-selective feature integration to clearly distinguish the tokens of the masked and unmasked region. (3) A novel inpainting region ID resampling technique to efficiently process videos of any length while maintaining ID coherence. By freezing the pre-trained context encoder and DiT backbone, and adding an ID-Adapter, we enhance the backbone’s attention sampling by concatenating the original key-value vectors with the inpainting region tokens. During inference, inpainting region tokens from previous clips are appended to the current key-value vectors, ensuring the long-term preservation of target IDs. Notably, our VideoPainter supports plug-and-play and user-customized control.
@@
  • We propose VideoPainter, the first dual-branch video inpainting framework that supports plug-and-play background controls.
  • We design a lightweight context encoder for efficient and dense background control, and inpainting region ID resampling for ID consistency in any-length video inpainting and editing.
@@
  • Experiments show VideoPainter achieves state-of-the-art performance across $8$ metrics including video quality, masked region preservation, and text alignment in video inpainting and editing.
2. Related Work
2.1. Video Inpainting
Video inpainting approaches can be broadly classified into two categories based on whether they possess generative capabilities:
@@ 2.1. Video Inpainting
These methods (Zhou et al., 2023; Li et al., 2022; Hu et al., 2020; Zhang et al., 2022a, b) leverage architecture priors to facilitate pixel propagation. This includes utilizing local perception of 3D CNNs (Chang et al., 2019a; Wang et al., 2019; Hu et al., 2020; Chang et al., 2019b), and exploiting the global perception of attention to retrieve and aggregate tokens with similar texture for filling masked video (Lee et al., 2019; Zeng et al., 2020; Liu et al., 2021; Zhang et al., 2022a). They also introduce various physical quantities, especially optical flow, as auxiliary conditions as it simplifies RGB pixel inpainting by completing less complex flow fields (Kim et al., 2019; Li et al., 2020; Zou et al., 2021; Xu et al., 2019; Gao et al., 2020; Zhang et al., 2022c, b). However, they are only effective for partial object occlusions with random masks but face significant limitations when inpaint fully masked regions due to insufficient contexts.
@@
Recent advances in generative foundation models (Rombach et al., 2022; Guo et al., 2023) have sparked numerous approaches that leverage additional modules or training strategies to extend backbones’ capabilities for video inpainting (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024). AVID (Zhang et al., 2024b) and COCOCO (Zi et al., 2024) represent the most related recent works. Both adopt a similar implementation by augmenting Stable Diffusion Inpainting (Rombach and Esser, 2022) with trainable temporal attention layers. This architecture includes per-frame region filling based on the image inpainting backbone and temporal smoothing with temporal attention. Despite showing promising results for both random and segmentation masks due to their generative abilities, they struggle to balance background preservation and foreground generation with text caption (Ju et al., 2024; Li et al., 2024) within the single backbone. AVID also explores any-length video inpainting by smoothing latent at segment boundaries and using the middle frame as the ID reference. In contrast, VideoPainter is a dual-branch framework by decoupling video inpainting into foreground generation and background-guided preservation. It employs an efficient context encoder to guide any pre-trained DiT, facilitating plug-and-play control. Furthermore, VideoPainter also introduces a novel inpainting region ID resampling technique that enables ID consistency in any-length video inpainting.
Table 1. Comparison of video inpainting datasets. Our VPData is the largest video inpainting dataset to date, comprising over 390K high-quality clips with segmentation masks, video captions, and masked region descriptions.
@@
2.2. Video Inpainting Datasets
Recent advances in segmentation (Ravi et al., 2024) have created many video segmentation datasets (Perazzi et al., 2016; Xu et al., 2018; Hong et al., 2023; Ding et al., 2023; Tokmakov et al., 2023; Darkhalil et al., 2022). Among these, DAVIS (Perazzi et al., 2016) and YouTube-VOS (Xu et al., 2018) have become prominent benchmarks for video inpainting due to their high-quality masks and diverse object categories. However, the existing datasets face two primary limitations: (1) insufficient scale to meet the data requirements of generative models, and (2) the absence of crucial control conditions necessary for generating masked objects such as video captions. In contrast, as shown in Tab. 10, we developed a scalable dataset pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), and constructed the largest video inpainting dataset to date with over 390K clips, each annotated with segmentation masks and dense video captions.
Figure 3. Dataset Construction Pipeline. It consists of five pre-processing steps: collection, annotation, splitting, selection, and captioning.
3. Method
@@
Figure 4. Model overview. The upper figure shows the architecture of VideoPainter. The context encoder performs video inpainting based on concatenation of the noisy latent, downsampled masks, and masked video latent via VAE. Features extracted by the context encoder are integrated into the pre-trained DiT in a group-wise and token-selective manner, where two encoder layers modulate the first and second halves of the DiT, respectively, and only the background tokens will be integrated into the backbone to prevent information ambiguity. The lower figure illustrates the inpainting ID region resampling with the ID Resample Adapter. During training, tokens of the current masked region are concatenated to the KV vectors, enhancing ID preservation of the inpainting region. During inference, the ID tokens of the last clip are concatenated to the current KV vectors, maintaining ID consistency with the last clip by resampling.
3.1. VPData and VPBench Construction Pipeline
@@
Collection. We chose Videvo and Pexels  [^1] Videvo: https://www.videvo.net/, Pexels: https://www.pexels.com/ as our data sources. We finally obtained around $450K$ videos from these sources.
@@
  • We employ the Recognize Anything Model (Zhang et al., 2024a) for open-set video tagging to identify primary objects.
  • Based on the detected object tags, we utilize Grounding DINO (Liu et al., 2023) to detect bounding boxes for objects at fixed intervals.
  • These bounding boxes serve as prompts for SAM2 (Ravi et al., 2024), which generates high-quality mask segmentations.
  • Then we employ rigorous filtering criteria: inter-frame mask area variation $\Delta<20\%$ and frame coverage maintained between $30\%-70\%$ to ensure reliable segmentation masks quality.
Splitting. Scene transitions may occur while tracking the same object from different angles, causing disruptive view changes. We utilize PySceneDetect (Castellano, 2024) to identify scene transitions and subsequently partition the masks. Then we segmented the sequences into 10-second intervals and discarded short clips (¡ 6s).
Selection. We employ $3$ key criteria: (1) Aesthetic Quality, evaluated using the Laion-Aesthetic Score Predictor (Schuhmann et al., 2022); (2) Motion Strength, predicted by optical flow measurements using the RAFT(Teed and Deng, 2020); and (3) Content Safety, assessed via the Stable Diffusion Safety Checker (Rombach et al., 2022).
@@
3.2. Dual-branch Inpainting Control
We incorporate masked video features into the pre-trained diffusion transformer (DiT) via an efficient context encoder, to decouple the background context extraction and foreground generation. This encoder processes a concatenated input of noisy latent, masked video latent, and downsampled masks. Specifically, the noisy latent provides information about the current generation. The masked video latent, extracted via VAE, aligns with the pre-trained DiT’s latent distribution. We apply cubic interpolation to downsample masks, ensuring dimensional compatibility between masks and latents.
@@ 3.2. Dual-branch Inpainting Control
3.3. Target Region ID Resampling
While current DiTs show promise in handling temporal dynamics (Kuaishou, 2024; Bian et al., 2024a), they struggle to maintain smooth transitions and long-term identity consistency.
Smooth Transition. Following AVID (Zhang et al., 2024b), we employ overlapping generation and weighted average to maintain consistent transitions. Additionally, we utilize the last frame of the previous clip (before overlap) as the first frame of the current clip’s overlapping region to ensure visual appearance continuity.
Identity Consistency. To maintain identity consistency in the long video, we introduce an inpainting region ID resampling method, as shown in lower Fig. 4. During training, we freeze both the DiT and the context encoder. Then we add trainable ID-Resample Adapters into the frozen DiT (LoRA), enabling ID resampling functionality. Specifically, tokens from the current masked region, which contain the desired ID, are concatenated with the KV vectors, thereby enhancing ID preservation in the inpainting region through additional KV resampling. During inference, we prioritize maintaining ID consistency with the inpainting region tokens from the previous clip, as it represents the most temporally proximate generated result. Specifically, given current $Q^{v}_{i}$, $K^{v}_{i}$, and $V^{v}_{i}$, we concatenate tokens containing ID information ($K^{id}_{i}$ and $V^{id}_{i}$) to the current KV pairs (during training, these are tokens from the current inpainting region; during inference, from the previous clip’s inpainting region). This forms new KV-vectors $[K^{v}_{i},K^{id}_{i}]$ and $[V^{v}_{i},V^{id}_{i}]$ (where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation), enabling the model to sample necessary ID information and better maintain ID consistency.
3.4. Plug-and-Play Control